Tyrone is a utility maximizer. His income is $100, which he can spend on cafeteria meals and on notepads. Each meal costs $5, and each notepad costs $2. At these prices Tyrone chooses to buy 16 cafeteria meals and 10 notepads.

The price of notepads falls to $1; the price of cafeteria meals remains the same. What are the new values for the vertical intercept, horizontal intercept, and slope?

Vertical Intercept:

Horizontal intercept:

Slope:
Answer #1

Answer

Let Meal be represented by M and Notepad be represented by N.

Assuming M on horizontal axis and N on vertical axis.

Budget constraint is given by :

(Price of M) * M + (Price of N) * N = Income

Here Price of M = 5, Price of N = 1 and Income = 100

=> 5M + 1N = 100

Vertical intercept is that value where M = 0 => 5*0 + 1N = 100 => N = 100

Hence, New vertical intercept = 100

Horizontal intercept is that value where N = 0 => 5*M + 1*0 = 100 => M = 20

Hence, New horizontal intercept = 20.

Budget line is given by : 5M + 1N = 100

=> N = 100 - 5M => dN/dM = -5 --------------------New budget line slope.

Hence, New budget line slope = -5